OpenStudy (anonymous):

Can anyone help solve this crazy equation, please!? 20^3+20^2(20/3)-20(20/3)^2-(20/3)^3

OpenStudy (anonymous):

\[20^{3}+20^{2}(20/3)-20(20/3)^{2}-(20/3)^{3}\]

OpenStudy (dumbcow):

i thought maybe it could be factored but i don't think so just use order of operations and use a calculator :)

OpenStudy (anonymous):

\[[27(8000)+9(8000)-3(8000)-(8000)]/27\]

OpenStudy (anonymous):

you can factor 20^3 out of everything

OpenStudy (anonymous):

[(8000)(27+9-3-1)]/27

OpenStudy (anonymous):

[(8000)(32)]/27

OpenStudy (dumbcow):

9481.48148

OpenStudy (anonymous):

256000/27

OpenStudy (anonymous):

9481.481481...................

OpenStudy (anonymous):

ok how aare you guys getting these values? where did the 27, 9,3 and 1 come from?

OpenStudy (anonymous):

27 is common denominator

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

it looks like it's 3.

OpenStudy (anonymous):

20^3+20^2(20/3)-20(20/3)^2-(20/3)^3 1st term (20^3) = (20^3)/1 2nd term 20^2(20/3) = (20^3)/3 3rd term -20(20/3)^2 = -(20^3)/9 4th term -(20/3)^3 = -(20^3)/27 Now let (20/3) = x so you have (x/1) + (x/3) - (x/9) - (x/27) Hope it helps!

OpenStudy (anonymous):

Actually, your answer turned on the light bulb, so to speak! I do understand, of course it is very confusing to me, but I get the 27, now I just need to take notes from your post up there and apply it to my work here. Thanks so much! That was great! Don't know what I would have done without the help here!